# Photometric classification of stars around the Milky Way’s central black hole

### v1

@@ Abstract
Methods. We use intermediate band (IB) photometry to identify candidates for massive young stars. To ensure robust classification, we apply three different, but complementary methods: Bayesian inference, a basic neural network, and a fast gradient-boosted trees algorithm.
Results. We obtained spectral energy distributions for 6590 stars, 1181 of which have been previously classified spectroscopically. We identify 351 stars that are classified as early types by all three classification methods, with 155 of them being newly identified candidates. The radial density profiles for late and early-type stars fit well with broken power laws, revealing a break radius of $9.2\pm 0.6^{\prime\prime}$ for early-type stars. The late-type stars show a core-like distribution around Sgr A* while the density of the early-type stars increases steeply towards the black hole, consistent with previous work. We infer a top-heavy IMF of the young stars near Sgr A* ($R<9^{\prime\prime}$), with a power-law of $1.6\pm 0.1$. At greater distances from Sgr A* a standard Salpeter/Kroupa IMF can explain the data. Additionally, we demonstrate that IB photometry can also constrain the metallicities of late-type stars, estimating metallicities for over 600 late-type stars.
@@ 1 Introduction
Over the last few decades, numerous studies have unveiled the presence of young (${\sim}2-8$ Myr old), massive stars within ${\sim}0.5$ pc of the black hole (e.g. Ghez et al., 2003; Paumard et al., 2006; Lu et al., 2008; Bartko et al., 2009; Genzel et al., 2010; Do et al., 2013; Yelda et al., 2014). This discovery is unexpected, as star formation is typically thought to be suppressed in the vicinity of a massive black hole due to its strong tidal field. There is a peculiar population of apparently main-sequence B stars within 0.08′′ of the black hole, called S-cluster. At a larger radius, the young stars are distributed in two main features: the clockwise (CW) rotating disk in the region ${\sim}1$′′ (${\sim}0.04$pc) to ${\sim}8$′′ (${\sim}0.32$) from the black hole, and the counterclockwise (CCW) disk at distances larger than $8^{\prime\prime}$ (${\sim}0.32$pc) from Sgr A*. However, the latter feature is controversial, with some studies supporting its existence (Genzel et al., 2003; Paumard et al., 2006; Bartko et al., 2009) while others refute it (Lu et al., 2008; Yelda et al., 2014). While all studies find a top-heavy mass function for the young stars near Sgr A*, there are significant discrepancies as to its exact power-law slope, from a value of -0.45 reported by Bartko et al. (2010) to -1.7 reported by Lu et al. (2013). This inconsistency may arise from the limited and different completeness of the data in these studies. Recently, von Fellenberg et al. (2022) examined an extensive area (${\sim}30^{\prime\prime}\times 30^{\prime\prime}$), the most extensive study of young stars in the GC to this date. They confirm the presence of the warped CW disk, an outer kinematic feature (F2), and the CCW feature at larger distances. Additionally, a previously unreported feature (F3) was identified beyond a projected radius of $R=8^{\prime\prime}$ from Sgr A*. The stars from the warped CW disk exhibit a top-heavy mass function, in contrast to the stars at larger distances.
An important question regards the formation mechanism of the young stars. While most studies support in-situ star formation from a massive gaseous disk, none can provide a complete explanation. Identifying more early-type stars will enhance our comprehension of their distribution and, consequently, shed light on their initial mass function (IMF), potentially providing further insights.
@@ 3.3.2 Local calibration
The calibration within the central region (i2, j2 in Fig. 26 in Appendix C) exhibits minimal sensitivity to systematic errors and statistical uncertainties because of the large number of exposures, closeness to the AO guide star and high density of calibration stars. We selected $29$ RC stars classified by Gautam et al. (2019) in this region and derived their mean SED, which was subsequently employed as our SED template for the next step (see Fig. 3).
@@ 3.4 CO band
$f(x)=a+b\cdot x+10^{-0.4\cdot c}\cdot e^{x\cdot d}$ (1)
@@ 3.5 Cross matching with spectroscopic studies
As a final step, we compared the spectroscopic classifications with ours used in Section 3.3, which led to the inclusion of a few more stars but also revealed some discrepancies. We re-examined the early and late stars from Feldmeier-Krause et al. (2017a), which have been assigned to different catalogue stars using different methods. The correct match was always identifiable for early-type stars. However, for late-type stars, in some cases, the Feldmeier-Krause et al. (2017a) classification likely resulted from two similar bright stars merging in the seeing-limited spectrum. In these cases, neither star was definitively assigned a certain type, and they are excluded from both the early and late samples in the following sections. Further details on interesting stars can be found in Appendix E. Figure 5 shows the CBD diagram for all spectroscopically classified stars: 982 late-type stars (red points) and 212 early-type stars (blue points).
We explored also some interesting stars in Appendix E.
@@ 4 Bayesian inference
After the initial approximate classification, we employed a more refined Bayesian analysis. We fitted models to the two branches in the CBD diagram (CBDD) corresponding to late and early-type stars (see Fig. 9). Subsequently, we inferred the probabilities of each star belonging to either the early-type or late-type branch. Throughout the following sections, we refer to the parts of the CBDD that predominantly contain early or late-type stars as the early-type branch and late-type branch, respectively. The prior is determined by the initially estimated density of spectroscopically classified GC stars in the respective branches.
@@ 4.2 Iteration on priors
$f(\rm{x})=[a\cdot e^{b\,x}+h\cdot e^{-(\rm{x}-\mu)^{2}/\rm{\sigma}^{2}}]\cdot[1-1/(1+e^{-c\cdot(x-d)})]$ (2)
@@
$p_{\textit{e}}=\frac{d(\textit{e},\textit{IB})\cdot d(\textit{e},\textit{r})}{<d(\textit{e},\textit{IB})>}$ (3)
@@
$p_{\textit{l}}=\frac{d(\textit{l},\textit{IB})\cdot d(\textit{l},\textit{r})}{<d(\textit{l},\textit{IB})>}$ (4)
@@
$p_{\textit{e-norm}}=\frac{p_{\textit{e}}}{p_{\textit{e}}+p_{\textit{l}}}$ (5)
This process is repeated multiple times until all changes become very small, typically around 36 iterations.
@@ 5.1 Multi-Layer Perceptron
Our MLP model is constructed with the following architecture: an input layer comprising 5 neurons with Exponential Linear Unit (ELU) activation, followed by a dropout layer with a 50% dropout rate. The subsequent hidden layer consists of 5 neurons using Rectified Linear Unit (ReLU) activation. The final layer, designed for binary classification distinguishing late- and early-type stars, comprises 1 neuron with a Sigmoid activation function. The model incorporates 66 trainable parameters. For optimization, the Adam (Adaptive Moment Estimation) optimizer is employed, utilizing a binary binary_crossentropy loss function designed for binary classification tasks. In our training configuration, we set the learning rate to 0.001, a critical parameter governing the magnitude of weight adjustments during training. To mitigate overfitting, we incorporated an early stopping callback with a patience of 4. During the model training process, we allocated 20% of the training data for validation purposes. The neural network completed 44 training epochs. The validation accuracy achieved 93.9%, with a false positive rate of 6.1% on the validation set. When evaluated on the testing dataset, the model demonstrated an accuracy of 95.8%, with a false positive rate of 4.2%.
@@ 6.1 Early-type and late-type star candidates
Figure 13 illustrates the positions of young candidates brighter than 16 magnitudes, detected using the MLP (red circles), Bayesian (green circles), and XGBoost method (violet circles).
We discuss some especially interesting candidates in Appendix E.
Figure 13: The positions of spectroscopically classified stars (blue points) and early-type candidates brighter than 16 magnitudes for the different methods have been superimposed on the image from 12 June 2004. The three distinct regions, as identified in von Fellenberg et al. (2022), are delineated by circles in the images. The right hand panel shows a zoom into the central 20′′, marked by the black square on the left.
@@ 6.2 Structure of the cluster
For studying the radial profile, we exclude all stars fainter than 16 magnitudes and those situated in regions where only two observed frames are available for a given epoch (specifically, the four corners), as illustrated in Fig. 16.
@@ 7.1 Distribution of the stars in the Galactic Centre
Figure 17: The CBD diagram illustrates stars selected within rings around Sgr A*, highlighting the diminishing influence of the young bright branch as the distance from the black hole increases.
@@
We now compare our results to the literature. Our IB photometry results cannot be compared to Buchholz et al. (2009) as their data was not published, and Plewa (2018) only provided uncalibrated data without uncertainties. It is worth noting that our study, along with the latter, utilizes a probabilistic approach in assigning probabilities of early type to stars, unlike the former classification-based method. Plewa (2018) identified 363 early-type candidates ($p>0.5$). We successfully matched 2756 of their 3165 stars with our catalogue, including 303 of their early-type candidates. The number of spectroscopically classified stars is greater in our catalogue than in the earlier work of Plewa (2018) and we can therefore cross-check their results. 28 of their early-type candidates are spectroscopically classified as late-type, and 64 as early-type. As for their candidates without a spectroscopic match (211), the majority is not early according to our probabilities, with 74% in the case of our Bayesian method. The majority of the early stars from Plewa (2018) do not lie in the early-type region of our CBDD. We conclude that the differences between their study and ours are significantly larger than the differences between our 3 methods at these magnitudes. To assess whether the difference is random or caused by a systematic issue, we counted the number of early stars in four bins of magnitude and radius, as shown in Table 7. There are systematic differences visible, with more early stars detected, particularly noticeable for faint and outer stars, compared to our three methods, all of which show close agreement. Additionally, there is a blue clump that is more prominent than in all of our datasets. This indicates suboptimal analysis, especially for faint stars. Another reason for the difference is that Plewa (2018) used only the early stars from Paumard et al. (2006), which included very few fainter early stars, making extrapolation more important in his analysis. We believe that the main reason for these differences lies in the way the data were reduced and analysed. We have set serious emphasis on careful photometry with robust uncertainty estimation as well as on exploring different ways of analysing the data.
@@
Buchholz et al. (2009) identified 312 new early-type candidates in addition to the 90 early-type stars that were spectroscopically identified at the time. We successfully matched 3908 of their 5914 stars with our catalogue, including 280 of their early-type classifications. We cross-checked their classifications with our compilation of spectroscopic data. 19 of their early-type candidates are spectroscopically classified as late-type, 139 as early-type. As for their 124 candidates without a spectroscopic match, the majority is not early according to our probabilities, with 59% in the case of our Bayesian method. This indicates better performance compared to Plewa (2018), aligning more closely with our findings. That is also confirmed when comparing the total number of early stars, as shown in Table 8. In most bins, the Buchholz et al. (2009) sum is between our 3 variants, except for faint stars further out - the most challenging region with the fewest stars. It is also possible that some of our early probabilities are unclassified by their work. Thus, this time, the main contributing factor is random differences in the CBD values. Similar to the case in Plewa (2018), the agreement is minimal for most RC stars and some brighter stars located beyond approximately $12^{\prime\prime}$. The increased depth also results in agreement for some stars fainter than the RC within about $7^{\prime\prime}$
@@ 7.2 K-band Luminosity Function
We determined a slope parameter for the KLF of late-type stars, obtaining a value of 0.25. This parameter exhibits only minor variations across different classification methods and radial ranges. Compared with previous studies, Buchholz et al. (2009) reported a steeper slope of 0.31$\pm$0.01 using stars up to $m=14.5$ mag, while Plewa (2018) obtained a steeper slope of 0.36 using stars up to 14.5 mag. The differences in slopes could be attributed to completeness corrections or the broader magnitude range considered in the current study. Notably, in the case of Plewa (2018), their RC is rather faint, which means that the bright end tail matters more for their fit, likely contributing to the observed steep slope. That observation is supported by the steeper slope, ranging from 0.29 to 0.32, obtained when excluding stars within the RC or those with fainter magnitudes.
@@ 7.4 Temperature and metallicity of late-type stars
We begin our analysis by examining the effective temperature of late-type stars, which is more directly related to the CBD than metallicity.
@@
We exclude stars with a metallicity $[M/H]<-2$, which we do not consider reliable. Our analysis of 25 stars with data from both works revealed no significant temperature offset between them. In total, we analyzed 394 stars, 51 from Do et al. (2015) and the majority (318 stars) from Feldmeier-Krause et al. (2017a).
When plotting the temperatures against CBD (Fig. 18) we find a linear relationship. This allows us to connect these two parameters via the following relationship:
$T_{\text{eff}}=(4157\pm 44)+(-1109\pm 101)\cdot\text{CBD}$ (6)
@@
$\begin{split}CBD=&1.08167498-0.00018242\cdot T_{\text{eff}}\\ &+0.07098599\cdot\text{[M/H]}-0.05281159\cdot\text{[M/H]}^{2}\end{split}$ (7)
@@
Next, we used the obtained relationship in Equation 7 to look into the metallicity. From the same 25 stars with metallicity as before, we found a consistent offset of 0.34 dex between the metallicities estimated by Do et al. (2015) and Feldmeier-Krause et al. (2017a). To align both datasets on the same stars, we adjusted by adding/subtracting half of this offset.
@@
Figure 19: Top: CBDD illustrating late stars with spectroscopic metallicity from Do et al. (2015) and Feldmeier-Krause et al. (2017a). We used 12 Gyrs Giant branch isochrones with different metallicities from Bressan et al. (2012); Chen et al. (2015); Marigo et al. (2017) to convert Teff into CBD values. Bottom: CBDD diagram of stars lacking spectroscopic metallicities and not classified as spectroscopic early. Metallicities are assigned using these isochrones. Both plots utilize the same colour scale and are restricted to stars brighter than 14.7, as this is where the presence of the RC, not included in our model, becomes significant. Stars encircled in grey represent those classified as likely early by the Bayesian method.
The fraction of low-metallicity stars determined from our photometric data aligns with spectroscopic assessments suggesting that warm/low-metallicity stars comprise up to 10% of the total population, as reported in studies such as Pfuhl et al. (2011); Do et al. (2015); Feldmeier-Krause et al. (2017a). While our estimate is in line with these findings, it remains uncertain due to the potential inclusion of early-type stars in our sample, as well as the likelihood of overlooking low-metallicity late-type stars, which may be misclassified as early-type due to weaker CO band heads, see also Appendix K. Addressing both factors comprehensively would necessitate modelling the spatial and magnitude distributions of all three classes (late metal-rich, late metal-poor, early), a task that falls beyond the scope of this paper.

### v2

@@ Abstract
Methods. We used intermediate band (IB) photometry to identify candidates for massive young stars. To ensure robust classification, we applied three different, but complementary methods: Bayesian inference, a basic neural network, and a fast gradient-boosted trees algorithm.
Results. We obtain spectral energy distributions for 6590 stars, 1181 of which have been previously classified spectroscopically. We identify 351 stars that are classified as early types by all three classification methods, with 155 of them being newly identified candidates. The radial density profiles for late and early-type stars fit well with broken power laws, revealing a break radius of $9.2\pm 0.6^{\prime\prime}$ for early-type stars. The late-type stars show a core-like distribution around Sgr A* while the density of the early-type stars increases steeply towards the black hole, consistent with previous work. We infer a top-heavy IMF of the young stars near Sgr A* ($R<9^{\prime\prime}$), with a power-law of $1.6\pm 0.1$. At greater distances from Sgr A* a standard Salpeter/Kroupa IMF can explain the data. Additionally, we demonstrate that IB photometry can also constrain the metallicities of late-type stars, estimating metallicities for over 600 late-type stars.
@@ 1 Introduction
Over the last few decades, numerous studies have unveiled the presence of young (${\sim}2-8$ Myr old), massive stars within ${\sim}0.5$ pc of the black hole (e.g. Ghez et al., 2003; Paumard et al., 2006; Lu et al., 2008; Bartko et al., 2009; Genzel et al., 2010; Do et al., 2013; Yelda et al., 2014). This discovery is unexpected, as star formation is typically thought to be suppressed in the vicinity of a massive black hole due to its strong tidal field. There is a peculiar population of apparently main-sequence B stars within 0.08′′ of the black hole, called ‘S-cluster’. At a larger radius, the young stars are distributed in two main features: the clockwise (CW) rotating disk in the region ${\sim}1$′′ (${\sim}0.04$pc) to ${\sim}8$′′ (${\sim}0.32$) from the black hole, and the counterclockwise (CCW) disk at distances larger than $8^{\prime\prime}$ (${\sim}0.32$pc) from Sgr A*. However, the latter feature is controversial, with some studies supporting its existence (Genzel et al., 2003; Paumard et al., 2006; Bartko et al., 2009) while others refute it (Lu et al., 2008; Yelda et al., 2014). While all studies find a top-heavy mass function for the young stars near Sgr A*, there are significant discrepancies as to its exact power-law slope, from a value of -0.45 reported by Bartko et al. (2010) to -1.7 reported by Lu et al. (2013). This inconsistency may arise from the limited and different completeness of the data in these studies. Recently, von Fellenberg et al. (2022) examined an extensive area (${\sim}30^{\prime\prime}\times 30^{\prime\prime}$), the most extensive study of young stars in the GC to this date. They confirm the presence of the warped CW disk, an outer kinematic feature (F2), and the CCW feature at larger distances. Additionally, a previously unreported feature (F3) was identified beyond a projected radius of $R=8^{\prime\prime}$ from Sgr A*. The stars from the warped CW disk exhibit a top-heavy mass function, in contrast to the stars at larger distances.
An important question regards the formation mechanism of the young stars. While most studies support in situ star formation from a massive gaseous disk, none can provide a complete explanation. Identifying more early-type stars will enhance our comprehension of their distribution and, consequently, shed light on their initial mass function (IMF), potentially providing further insights.
@@ 3.3.2 Local calibration
The calibration within the central region (i2, j2 in Fig. 26) exhibits minimal sensitivity to systematic errors and statistical uncertainties because of the large number of exposures, closeness to the AO guide star and high density of calibration stars. We selected $29$ RC stars classified by Gautam et al. (2019) in this region and derived their mean SED, which was subsequently employed as our SED template for the next step (see Fig. 3).
@@ 3.4 CO band
$f(x)=a+b\cdot x+10^{-0.4\cdot c}\cdot e^{x\cdot d},$ (1)
@@ 3.5 Cross matching with spectroscopic studies
As a final step, we compared the spectroscopic classifications with ours used in Section 3.3, which led to the inclusion of a few more stars but also revealed some discrepancies. We re-examined the early and late stars from Feldmeier-Krause et al. (2017a), which have been assigned to different catalogue stars using different methods. The correct match was always identifiable for early-type stars. However, for late-type stars, in some cases, the Feldmeier-Krause et al. (2017a) classification likely resulted from two similar bright stars merging in the seeing-limited spectrum. In these cases, neither star was definitively assigned a certain type, and they are excluded from both the early and late samples in the following sections. Further details on interesting stars can be found in Appendix E. Figure 5 shows the CBD diagram for all spectroscopically classified stars: 982 late-type stars (red points) and 212 early-type stars (blue points). We explored also some interesting stars in Appendix E.
@@ 4 Bayesian inference
After the initial approximate classification, we employed a more refined Bayesian analysis. We fitted models to the two branches in the CBD diagram (CBDD) corresponding to late and early-type stars (see Fig. 9). Subsequently, we inferred the probabilities of each star belonging to either the early-type or late-type branch. Throughout the following sections, we refer to the parts of the CBDD that predominantly contain early or late-type stars as the ‘early-type branch’ and ‘late-type branch’, respectively. The prior is determined by the initially estimated density of spectroscopically classified GC stars in the respective branches.
@@ 4.2 Iteration on priors
$f(\rm{x})=[a\cdot e^{b\,x}+h\cdot e^{-(\rm{x}-\mu)^{2}/\rm{\sigma}^{2}}]\cdot[1-1/(1+e^{-c\cdot(x-d)})],$ (2)
@@
$p_{\textit{e}}=\frac{d(\textit{e},\textit{IB})\cdot d(\textit{e},\textit{r})}{<d(\textit{e},\textit{IB})>}.$ (3)
@@
$p_{\textit{l}}=\frac{d(\textit{l},\textit{IB})\cdot d(\textit{l},\textit{r})}{<d(\textit{l},\textit{IB})>}.$ (4)
@@
$p_{\textit{e-norm}}=\frac{p_{\textit{e}}}{p_{\textit{e}}+p_{\textit{l}}}.$ (5)
This process was repeated multiple times until all changes become very small, typically around 36 iterations.
@@ 5.1 Multi-layer perceptron
Our MLP model was constructed with the following architecture: an input layer comprising 5 neurons with ‘Exponential Linear Unit’ (ELU) activation, followed by a dropout layer with a 50% dropout rate. The subsequent hidden layer consists of 5 neurons using ‘Rectified Linear Unit’ (ReLU) activation. The final layer, designed for binary classification distinguishing late- and early-type stars, comprises 1 neuron with a ‘Sigmoid’ activation function. The model incorporates 66 trainable parameters. For optimization, the ‘Adam’ (Adaptive Moment Estimation) optimizer is employed, utilizing a binary ‘binary_crossentropy’ loss function designed for binary classification tasks. In our training configuration, we set the learning rate to 0.001, a critical parameter governing the magnitude of weight adjustments during training. To mitigate overfitting, we incorporated an early stopping callback with a patience of 4. During the model training process, we allocated 20% of the training data for validation purposes. The neural network completed 44 training epochs. The validation accuracy achieved 93.9%, with a false positive rate of 6.1% on the validation set. When evaluated on the testing dataset, the model demonstrated an accuracy of 95.8%, with a false positive rate of 4.2%.
@@ 6.1 Early-type and late-type star candidates
Figure 13 illustrates the positions of young candidates brighter than 16 magnitudes, detected using the MLP (red circles), Bayesian (green circles), and XGBoost method (violet circles). We discuss some especially interesting candidates in Appendix E.
Figure 13: Map illustrating the positions of spectroscopically classified stars (blue points) and early-type candidates brighter than 16 magnitudes for the different methods, superimposed on the image from 12 June 2004. The three distinct regions, as identified in von Fellenberg et al. (2022), are delineated by circles in the images. The right hand panel shows a zoom into the central 20′′, marked by the black square on the left.
@@ 6.2 Structure of the cluster
For studying the radial profile, we excluded all stars fainter than 16 magnitudes and those situated in regions where only two observed frames are available for a given epoch (specifically, the four corners), as illustrated in Fig. 16.
@@ 7.1 Distribution of the stars in the Galactic Centre
Figure 17: CBD diagrams computed for stars in different rings around Sgr A* highlight the diminishing influence of the young bright branch as the distance from the black hole increases.
@@
We now compare our results to the literature. Our IB photometry results cannot be compared to Buchholz et al. (2009) as their data was not published, and Plewa (2018) only provided uncalibrated data without uncertainties. It is worth noting that our study, along with the latter, utilizes a probabilistic approach in assigning probabilities of early type to stars, unlike the former classification-based method. Plewa (2018) identify 363 early-type candidates ($p>0.5$). We successfully matched 2756 of their 3165 stars with our catalogue, including 303 of their early-type candidates. The number of spectroscopically classified stars is greater in our catalogue than in the earlier work of Plewa (2018) and we can therefore cross-check their results. 28 of their early-type candidates are spectroscopically classified as late-type, and 64 as early-type. As for their candidates without a spectroscopic match (211), the majority is not early according to our probabilities, with 74% in the case of our Bayesian method. The majority of the early stars from Plewa (2018) do not lie in the early-type region of our CBDD. We conclude that the differences between their study and ours are significantly larger than the differences between our 3 methods at these magnitudes. To assess whether the difference is random or caused by a systematic issue, we counted the number of early stars in four bins of magnitude and radius, as shown in Table 7. There are systematic differences visible, with more early stars detected, particularly noticeable for faint and outer stars, compared to our three methods, all of which show close agreement. Additionally, there is a blue clump that is more prominent than in all of our datasets. This indicates suboptimal analysis, especially for faint stars. Another reason for the difference is that Plewa (2018) used only the early stars from Paumard et al. (2006), which included very few fainter early stars, making extrapolation more important in his analysis. We believe that the main reason for these differences lies in the way the data were reduced and analysed. We have set serious emphasis on careful photometry with robust uncertainty estimation as well as on exploring different ways of analysing the data.
@@
Buchholz et al. (2009) identify 312 new early-type candidates in addition to the 90 early-type stars that were spectroscopically identified at the time. We successfully matched 3908 of their 5914 stars with our catalogue, including 280 of their early-type classifications. We cross-checked their classifications with our compilation of spectroscopic data. 19 of their early-type candidates are spectroscopically classified as late-type, 139 as early-type. As for their 124 candidates without a spectroscopic match, the majority is not early according to our probabilities, with 59% in the case of our Bayesian method. This indicates better performance compared to Plewa (2018), aligning more closely with our findings. That is also confirmed when comparing the total number of early stars, as shown in Table 8. In most bins, the Buchholz et al. (2009) sum is between our 3 variants, except for faint stars further out - the most challenging region with the fewest stars. It is also possible that some of our early probabilities are unclassified by their work. Thus, this time, the main contributing factor is random differences in the CBD values. Similar to the case in Plewa (2018), the agreement is minimal for most RC stars and some brighter stars located beyond approximately $12^{\prime\prime}$. The increased depth also results in agreement for some stars fainter than the RC within about $7^{\prime\prime}$
@@ 7.2 K-band luminosity function
We determine a slope parameter for the KLF of late-type stars, obtaining a value of 0.25. This parameter exhibits only minor variations across different classification methods and radial ranges. Compared with previous studies, Buchholz et al. (2009) reported a steeper slope of 0.31$\pm$0.01 using stars up to $m=14.5$ mag, while Plewa (2018) obtained a steeper slope of 0.36 using stars up to 14.5 mag. The differences in slopes can be attributed to completeness corrections or the broader magnitude range considered in the current study. Notably, in the case of Plewa (2018), their RC is rather faint, which means that the bright end tail matters more for their fit, likely contributing to the observed steep slope. That observation is supported by the steeper slope, ranging from 0.29 to 0.32, obtained when excluding stars within the RC or those with fainter magnitudes.
@@ 7.4 Temperature and metallicity of late-type stars
We began our analysis by examining the effective temperature of late-type stars, which is more directly related to the CBD than metallicity.
@@
We excluded stars with a metallicity $[M/H]<-2$, which we do not consider reliable. Our analysis of 25 stars with data from both works revealed no significant temperature offset between them. In total, we analyzed 394 stars, 51 from Do et al. (2015) and the majority (318 stars) from Feldmeier-Krause et al. (2017a).
When plotting the temperatures against CBD (Fig. 18) we found a linear relationship. This allows us to connect these two parameters via the following relationship:
$T_{\text{eff}}=(4157\pm 44)+(-1109\pm 101)\cdot\text{CBD},$ (6)
@@
$\begin{split}CBD=&1.08167498-0.00018242\cdot T_{\text{eff}}\\ &+0.07098599\cdot\text{[M/H]}-0.05281159\cdot\text{[M/H]}^{2}.\end{split}$ (7)
@@
Next, we used the obtained relationship in Equation 7 to look into the metallicity. From the same 25 stars with metallicity as before, we found a consistent offset of 0.34 dex between the metallicities estimated by Do et al. (2015) and Feldmeier-Krause et al. (2017a). To align both datasets on the same stars, we adjusted by adding or subtracting half of this offset.
@@
Figure 19: Metallicity estimation analysis. The top panel illustrates the CBDD of late stars with spectroscopic metallicity from Do et al. (2015) and Feldmeier-Krause et al. (2017a). We used 12 Gyrs Giant branch isochrones with different metallicities from Bressan et al. (2012); Chen et al. (2015); Marigo et al. (2017) to convert Teff into CBD values. The bottom panel shows the CBDD diagram of stars lacking spectroscopic metallicities and not classified as spectroscopic early. Metallicities are assigned using these isochrones. Both plots utilize the same colour scale and are restricted to stars brighter than 14.7, as this is where the presence of the RC, not included in our model, becomes significant. Stars encircled in grey represent those classified as likely early by the Bayesian method.
The fraction of low-metallicity stars determined from our photometric data aligns with spectroscopic assessments suggesting that warm,low-metallicity stars comprise up to 10% of the total population, as reported in studies such as Pfuhl et al. (2011); Do et al. (2015); Feldmeier-Krause et al. (2017a). While our estimate is in line with these findings, it remains uncertain due to the potential inclusion of early-type stars in our sample, as well as the likelihood of overlooking low-metallicity late-type stars, which may be misclassified as early-type due to weaker CO band heads, see also Appendix K. Addressing both factors comprehensively would necessitate modelling the spatial and magnitude distributions of all three classes (late metal-rich, late metal-poor, early), a task that falls beyond the scope of this paper.